OpenStudy (anonymous):

What is the line of symmetry for the graph of y = -3x2 + 12x - 11

OpenStudy (anonymous):

y = -3x^2 + 12x - 11 the x-coordinate of the vertex is -b/2a = -12/(-2*3) = 2. x= 2 is the line of symmetry

OpenStudy (anonymous):

Or via derivative, -6x +12 = 0, x = 2.

OpenStudy (anonymous):

Aeterna, did you have a question?